OpenStudy (anonymous):

Find The Surface Area of The Triangular Prism 1. 81 in. 2 2. 70.8 in.2 3. 63in.2 4. 78.6 in2 (i picked 63in.2)

jigglypuff314 (jigglypuff314):

Could you attach the picture of the prism? or use the "Draw" tool below? :)

OpenStudy (anonymous):

@jigglypuff314 @mathstudent55 @geekfromthefutur @

OpenStudy (anonymous):

yeah got it :)

jigglypuff314 (jigglypuff314):

alright :) so you found that 3*7 is one of the rectangular sides then there's three of them so you did 3*7*3 = 63 right? :) but don't forget that you have two other triangles to add to it area of triangle = (1/2)(base)(height) = (1/2) (3) (2.6) and there are two triangles

OpenStudy (anonymous):

i looked up find the area of it and tried to get everything in but did it wrong i still got 63 though

OpenStudy (geekfromthefutur):

Did u get it

jigglypuff314 (jigglypuff314):

hmmm... You should add 2(1/2)(3)(2.6) to the 63 because you want to account for the two end triangles :)

OpenStudy (anonymous):

2(1/2)(3)(2.6) + 63 = 70.8

jigglypuff314 (jigglypuff314):

Correct :)

OpenStudy (anonymous):

yayyy :)

OpenStudy (anonymous):

thanks ;D

jigglypuff314 (jigglypuff314):

You're Welcome ^_^

OpenStudy (anonymous):

Medal given ^^

jigglypuff314 (jigglypuff314):

Thanks :)

OpenStudy (anonymous):

np ;D